OpenStudy (anonymous):

Write 3 Ordered Pairs 2y=-3x+2

OpenStudy (anonymous):

(0, 1) | 2(1) = -3(0) + 2 (2, 2) | 2(2) = -3(2) + 2 (4, -5) | 2(-5) = -3(4) + 2

OpenStudy (anonymous):

(-5, 4) for the last one